A ball player catches a ball 3.79 s after throwing it vertically upward. 


 Part A With what speed did he throw it?

 Part B What height did it reach?  
Answer #1

total time of flight , t = 3.79 s

part A)

speed of the ball = g * t/2

speed of the ball = 9.8 * 3.79/2

speed of the ball = 18.6 m/s

the speed of the ball is 18.6 m/s

part B)

mximum height reached = initial speed^2/(2g)

mximum height reached = 18.6^2/(2 * 9.8)

mximum height reached = 17.6 m